Could someone please help me solve this question: Solve the following system of two linear equations by substitution: x + 2 y – 8 =0 & 2 x + 4 y + 4 =0.
Noting each equation represents a line, draw both lines and interpret the results.
Thank you, for taking the time to help me.

Answer by stanbon(48545)

Solve the following system of two linear equations by substitution: x + 2 y – 8 =0 & 2 x + 4 y + 4 =0.
Noting each equation represents a line, draw both lines and interpret the results.
-----------------------
1st: x + 2 y = 8
2nd: 2 x + 4 y = - 4
-------------------
Multiply the 1st be 2 to get:
3rd: 2x + 4y = 16
-----------
Subtract 2nd from 3rd to get:
0 = 20
That is a contradiction that happened because we assumed the
"x" in one equation was the same as the "x" in the other, and
the same for the "y".
They are not the same; the system in called "inconsistent"
------------------
If you were to graph the two equations you would get two
parallel lines.
there is no solution for the system of equations.
==================
Cheers,
Stan H.